OpenStudy (anonymous):

Inverse functions w/ logarithms. If g(x) = 3 + x + e^x, find g-1(4) I know that to set it up we do something akin to g(x) = 3 + x + e^x y = 3 + x + e^x x + e^x = y - 3 here is where I'm confused on what to do. do I take the natural log of everywhere, giving me, ln x + x = ln (y-3)? I don't know how to single out the x's.

OpenStudy (anonymous):

g(x)=y=3+x+(e^x) in order to get inverse of y interchange x and y so we have x=3+y+e^y => y+e^y=x-3 eqn 1 from above we can get get g-1(x) as function of x now we have to find g-1(4) so place x=4 in eqn 1 we get y+e^y=1 we know that e^0=1 so g-1(4)=0 Ans I hope the solution is explained properly